OpenStudy (anonymous):

What integer needs to be added to the expression to create a perfect square? x^2 - 2x

OpenStudy (anonymous):

(a-b)^2 = a^2 - 2*a*b +b^2 so here we have (x- ...)^2 =x^2 -2*x*1+ .... i hope you can understand that b-->1